# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

### v2

@@ 6 Conclusion
This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem (in both file-path identification task and function reproduction task), and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points for file-path identification task) despite their popularity and likely training data inclusion. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
These findings suggest that the improvements in SWE-Bench Verified performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.
Appendix A Appendix
A.1 Full Prompt Template
@@ DISCUSSION
RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: from sklearn import linear_model import numpy as np x = np.random.randn(100, 30) y = np.random.normal(size=100) model = linear_model.RidgeClassifierCV( alphas=np.arange(0.1, 1000, 0.1), normalize=True, store_cv_values=True ).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
@@
A.2 Additional Results
[Additional tables and figures supporting the main findings]
Model Mentioned Not Mentioned Overall
Single GT Multi GT Overall Single GT Multi GT Overall Single GT Multi GT Overall
claude-3-5-sonnet 88.60 71.43 85.93 51.90 44.90 50.96 61.63 52.86 60.40
claude-3-7-sonnet 92.11 85.71 91.11 60.13 57.14 59.73 68.60 65.71 68.20
claude-sonnet-4 91.23 85.71 90.37 65.51 53.06 63.84 72.33 62.86 71.00
claude-opus-4 92.11 95.24 92.59 67.41 61.22 66.58 73.95 71.43 73.60
gpt-4o (May) 91.23 85.71 90.37 51.27 55.10 51.78 61.86 64.29 62.20
gpt-4o 88.60 80.95 87.41 53.16 51.02 52.88 62.56 60.00 62.20
gpt-4.1 85.09 71.43 82.96 56.65 55.10 56.44 64.19 60.00 63.60
o3-mini 91.23 85.71 90.37 60.13 63.27 60.55 68.37 70.00 68.60
o3 92.98 76.19 90.37 71.84 65.31 70.96 77.44 68.57 76.20
o4-mini 92.98 80.95 91.11 59.49 48.98 58.08 68.37 58.57 67.00
Table 3: Accuracy of model-predicted file paths on SWE-Bench Verified, stratified by whether any GT path is explicitly mentioned in the problem description. Results are further broken down into three categories: Single GT Path (tasks where the GT patch modifies only a single execution path), Multiple GT Paths (tasks involving modifications across multiple execution paths), and Overall (all tasks). Among the task instances, 135 mention at least one GT path, while 365 do not.

### v3

@@ 6 Conclusion
This work presents evidence that current evaluations of LLM coding abilities may be compromised by benchmark contamination and memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific, evidenced by graduated performance decay within SWE-Bench ecosystem (in both file-path identification and function reproduction tasks), and repository-bias memorization, demonstrated by substantial performance drops on external repos (up to 47 percentage points for file-path identification) despite their popularity. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns of SWE-Bench Verified’s data.
These findings suggest that the improvements in SWE-Bench Verified performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks with temporal controls to prevent training data contamination, cross-repository validation to test generalization of these models beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than overfitting on dataset-specific artifacts.
Appendix A SWE-Bench C# Details
We show the repository distribution of SWE-Bench C# in Figure 6. This benchmark is useful for tuning agents problem-solving approach for bugfixing, feature addition, and refactoring in the C# and .NET ecosystem.
Figure 6: Distribution of repositories in SWE-Bench C#.
Appendix B Prompt Templates and Examples
B.1 File-Path Identification Prompt Template
We present the prompt template for the File-Path Identification task in Figure 7. The prompt begins with the repository name, followed by the issue description, which serves as the problem statement for SWE-Bench Verified tasks, and concludes with instructions for the model to reason through the problem and output a single predicted buggy file path.
B.2 SWE-Bench Verified Task Example
Figure 8 is a reduced example for SWE-Bench Verified instance with ID: scikit-learn__scikit-learn-10297. In practice, problem statements similar to this will be provided to LLMs in the file-path identification task, and the function reproduction task as a part of the prompt.
@@ DISCUSSION
RidgeClassifierCV store_cv_values Issue Error: TypeError: __init__() got an unexpected keyword argument ’store_cv_values’ Repro: 1from sklearn import linear_model 2import numpy as np 3 4x = np.random.randn(100, 30) 5y = np.random.normal(size=100) 6 7model = linear_model.RidgeClassifierCV( 8 alphas=np.arange(0.1, 1000, 0.1), 9 normalize=True, 10 store_cv_values=True 11).fit(x, y) Note: store_cv_values is not a valid parameter, yet some attributes depend on it.
@@
B.3 SWE-Bench C# Example
Figure 9 shows the problem statement for instance: ardalis__cleanarchitecture-530 in SWE-Bench C#.
Issue: Request to Create New ToDoItems, instance: [ardalis__cleanarchitecture-530] 1Please show how you would create new ToDoItems 2<!-- Do Not Delete This! feature_request_template --> 3<!-- Please search existing issues to avoid creating duplicates. --> 4 5<!-- Describe the feature youd like. --> 6 7In this template, you aggregate ToDoItems within the ProjectAggregate. However, you also stop short of showing how you would, via API, author new ToDoItem instances. I think it would be beneficial to show how you would create these new child items of a project. 8 9I think it would be wonderful to see: 10- How you would define this in an ApiEndpoint. 11- How you would incorporate this within the ProjectAggregate. 12- How you would codify a Many-to-Many relationship as opposed to the One-To-Many relationship between Project and ToDoItem. 13 - Is this implied by the Project <-> Contributor relationship? 14 - What if the Contributor to a Project had a title within that Project? Senior Contributor vs Junior Contributor? If that were the case, what ApiEndpoint would the management of that be within your example domain? 15 16Thanks for such a great template!
Figure 9: Example issue description requesting creation of new ToDoItem instances
B.4 Function Reproduction Task Prompt Template
We present the prompt template for the Function Reproduction task in Figure 10. The prompt begins with the repository name, followed by hints indicating which functions need to be implemented, and only the names and the file paths are included in the hints. It then provides the complete file content, where the buggy functions are removed and replaced by comments marking the index of each missing function. Finally, the prompt instructs the model to reason through the task and output the reconstructed content for all buggy functions identified in the provided file.
@@ B.4 Function Reproduction Task Prompt Template
Appendix C Additional Experiment Results
C.1 Comparison of 5-Gram Overlap for Buggy V.S. Ground Truth Code Snippet
Figure 11: Difference in models’ 5-gram overlap ratio when comparing against the ground truth code snippet versus the buggy code snippet in the function reproduction task across benchmarks. A higher difference indicates that the model’s solution is more similar to the ground truth.
We also report the difference of 5 gram overlap
$\Delta_{5}=\text{overlap}_{5}(\hat{f},f_{\text{GT}})-\text{overlap}_{5}(\hat{f},f_{\text{buggy}}),$
where $\text{overlap}_{5}(x,y)$ is the 5 gram overlap of $x$ and $y$. Positive values indicate the model output more closely matches the patched implementation; negative values indicate closer similarity to the buggy code.
Figure 11 summarizes $\Delta_{5}$ across benchmarks and model families. Three broad patterns emerge.
(1) Provider‑level asymmetry on legacy benchmarks.
Anthropic’s Claude models show consistently positive differentials on SWE‑Bench Verified, Full-SWE-Bench, and Outside-Repo Tasks (roughly +2 to -6 percentage points), suggesting that these models retain verbatim shards of the patched implementations. OpenAI models yield smaller positives (generally less than +2pp) over the same sets, indicating weaker memorization of those specific fixes.
(2) Freshness stress test.
All models, except Claude 3.5 Sonnet and GPT 4.1, produce negative $\Delta_{5}$ on SWE‑Bench Extra. Because this split is mined largely from post‑cutoff commits, the patches are unlikely to be in training data. The systematic shift toward the buggy reference therefore supports our contamination‑mitigation claim: when the patched code is unseen, models do not memorize it from context.
(3) Language / ecosystem skew.
On the SWE C# subset, OpenAI’s GPT‑4* family exhibits the largest positive differentials (~+3 to 5pp), exceeding Claude. This may suggest heavier exposure to patched C# / .NET code in OpenAI training data (or lighter deduplication of that ecosystem) compared to Anthropic.
Taken together, the function reproduction experiment provides a direct, contrastive probe of code memorization: when historical (patched) code is in a model’s training mix, we observe measurable positive $\Delta_{5}$; when the code post‑dates the model (SWE‑Bench Extra), the signal disappears or reverses. Additionally, the high tendency towards the fixed version of code on SWE-Bench Verified and SWE-Bench C# indicates higher possibility of benchmark memorization on these two settings. These results underscore the necessity of freshness‑controlled benchmarks when interpreting coding model performance and memorization risk.